why mass of proton=1837*mass of electron


 

Grade:

2 Answers

B.V.Suhas Sudheendhra
12 years ago

Mass of an electron = 9.11 x 10-31kg

Mass of a proton = 1.66 x 10-27kg

therefore mass of proton= 1.66*10^-27=1837* 9.11*10^-31kg


Nimon B
12 years ago

It is all by measurements.